# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

### v1

@@ Abstract.
Decompiler is a specialized type of reverse engineering tool extensively employed in program analysis tasks, particularly in program comprehension and vulnerability detection. However, current Solidity smart contract decompilers face significant limitations in reconstructing the original source code. In particular, the bottleneck of SOTA decompilers lies in inaccurate method identification, incorrect variable type recovery, and missing contract attributes. These deficiencies hinder downstream tasks and understanding of the program logic. To address these challenges, we propose SmartHalo, a new framework that enhances decompiler output by combining static analysis (SA) and large language models (LLM). SmartHalo leverages the complementary strengths of SA’s accuracy in control and data flow analysis and LLM’s capability in semantic prediction. More specifically, SmartHalo constructs a new data structure - Dependency Graph (DG), to extract semantic dependencies via static analysis. Then, it takes DG to create prompts for LLM optimization. Finally, the correctness of LLM outputs is validated through symbolic execution and formal verification. Evaluation on a dataset consisting of 465 randomly selected smart contract methods shows that SmartHalo significantly improves the quality of the decompiled code, compared to SOTA decompilers (e.g., Gigahorse). Notably, integrating GPT-4o with SmartHalo further enhances its performance, achieving precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
Smart Contract, Decompilation, Static Analysis, Large Language Model
ccs: Software and its engineeringccs: Software and its engineering Software creation and managementccs: Software creation and management Software verification and validation
1. Introduction
@@
Decompiler is a specific type of reverse engineering tool widely used for program analysis, such as program comprehension and vulnerability detection (Dinesh et al., 2020; Liao et al., 2024). For smart contracts written in Solidity, the goal of the decompiler is to recover the machine-executable code (e.g., the EVM bytecode (EVM, [n. d.])) back to the original source code written by developers (e.g., the Solidity program (Sol, [n. d.])). Unfortunately, in most cases, decompilers are incapable of fully reconstructing the original developer-written code, as the compiler irreversibly subverts crucial information for optimization (Hu et al., 2024). In particular, in Solidity EVM, critical information such as method boundaries, identifier names, and types are absent from the decompiler output. These limitations bring a number of negative impacts to its downstream tasks, as pointed out by quite a lot of prior research (Hu et al., 2024; Liao et al., 2022; Lai and Luo, 2020).
@@
State-of-the-art solidity decompilers, such as Gigahorse (Grech et al., 2019) and Etherscan decompiler (eth, [n. d.]), mainly shares the following three deficiency in their output: (1) inaccurate method identification: current decompilers often fail to precisely determine method boundaries, leading to the omission of important methods or erroneous method range identification. For example, decompilers often incorrectly recover multiple methods as a single method, which obviously increases complexity in certain tasks (e.g., method-level similarity comparison). (2) inaccurate variable type recovery: decompilers may produce type errors that are inconsistent with static domain rules. For instance, the decompiler often ignores predefined type (e.g., the return value of hash function keccak256 is the type of bytes32), while recovering them as the type of uint256 uniformly. (3) lack of contract attributes. Smart contracts employ state variables to record critical contract attributes (e.g., Asset, identity, router). Although these contract attributes are explicitly stated in the source code through meaningful variable names and annotations, they are entirely omitted in the decompiled code. As highlighted in prior work (Liao et al., 2022), contract attribute is crucial for vulnerability detection tasks. In summary, the above deficiencies impede reverse engineers from understanding the program logic, and also adversely affect downstream tasks such as vulnerability detection and component analysis.
@@
To the best of our knowledge, there has been very limited research (Liao et al., 2022; Chen et al., 2022b; Zhao et al., 2023) aimed at optimizing deficiencies for smart contract decompiler output. These research are far from satisfied, due to their limited scope and approach-wise weaknesses. More specifically, SmartDagger (Liao et al., 2022), a framework for detecting cross-contract vulnerability, can partially recover contract attributes (e.g., asset) for state variables from the decompiled bytecode. SigRec (Chen et al., 2022b) and DeepInfer (Zhao et al., 2023) can partially recover already-known method signatures and their corresponding parameter types. However, SigRec (Chen et al., 2022b) is heavily relied on pre-defined heuristics according to EVM instructions. Besides, both SmartDagger and DeepInfer are based on deep-learning models (e.g., deep neural networks) for information recovery. Therefore, their capability are quite limited to the model-scale and training datasets. For example, they suffers performance degradation when confronted with newly-emerging or rare contracts that are not exist in the training datasets.
@@
Our work.  In this paper, we propose SmartHalo, a novel framework designed to optimize the output of existing smart contract decompilers. In particular, SmartHalo aims at recovering method boundaries, variable types, and contract attributes through a novel combination of static analysis (SA) and large language models (LLMs). The output of SmartHalo can facilitate a number of downstream tasks in program comprehension and security analysis. For example, eliminating false positives and false negatives for vulnerability detection.
@@
The key insight of our research is that SA and LLMs can collaboratively enhance the output of existing decompilers. Specifically, the advantage of SA lies in its soundness, as it is accurate in handling the optimization targets that are with complex static constraints (He et al., 2018). Consider the optimization of variable types as an example. The types of state variables (e.g., address) are inherently related to the program expressions that access or modify them (e.g., msg.sender). Therefore, SA is well-suited to analyze these intricate static constraints (e.g., predefined functions) and refine the types of state variable accordingly. In the meantime, the strength of LLM lies in its completeness, as it possesses the flexibility to predict the optimization targets that lack static constraints (Chen et al., 2022a). For instance, with the remarkable capability of few-shot in-context learning, LLMs are flexible to predicting the types for certain local variables, which are rarely encountered in the program and often incapable of inferring their types based on the types of other variables.
@@
SmartHalo extracts the static dependencies within the program through SA, and then leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these static dependencies. This process involves three main steps, including dependency-based semantic extraction (Section 4.1), LLM-driven semantic enrichment (Section 4.2), and correctness verification (Section 4.3). In the first step, SmartHalo extracts three types of dependencies from the decompiled code, including state dependency (e.g., read and write on state variable), control-flow dependency, and type dependency, to construct a fine-grained Dependency Graph (DG). The DG contains static domain knowledge (i.e., dependencies) necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract. In the second step, SmartHalo uses the DG to create prompts that include optimization target contexts (i.e., methods or variables), optimization output candidates (e.g., type or attribute), and chain-of-thoughts that represent inference steps of static analysis for optimization targets. By learning the prompts, LLMs enable the joint optimization on decompiler output with the advantages of both SA and LLMs.
@@
A common challenge in LLM-adaptation is to handling its potential hallucination. In our task, LLMs unexpectedly alter the program behaviors of the original decompiled code, and introduce inference errors that contradict common static domain rules. As countermeasure, SmartHalo conducts the rigorous correctness verification for the LLM output. Specifically, SmartHalo utilizes symbolic execution and formal verification to validate the program-behavior equivalence between the original decompiled code and optimized code. In addition, SmartHalo integrates a set of static rules to identify and reject inference errors that violate static domain knowledge. To this end, the erroneous LLM outputs are combined with the verification suggestion to re-transmit to LLM for recurring optimization, for further eliminating the errors.
@@
Evaluation.  To evaluate the effectiveness of SmartHalo, we randomly select 500 methods from the largest open-source contract dataset (xbl, [n. d.]), and manually labelled a dataset containing 456 pairs of source code and decompiler outputs of smart contract methods (44 methods encountered decompilation error). The evaluation results indicate that, compared to the original decompiler output (i.e., Gigahorse (Grech et al., 2019)), SmartHalo (with GPT-3.5) improves the precision of method boundary identification by 30.03% and enhances the precision of variable type inference by 12.01%. Additionally, SmartHalo successfully recovers 68.06% of contract attributes, significantly outperforming SmartDagger, which only recovers 10.07% of contract attributes. With the help of SmartHalo, 60.22% of optimized codes can directly be recompiled using the Solidity compiler, whereas the original decompiled output can not support recompilation. We also prove that the optimized output of SmartHalo enhances the effectiveness of downstream tasks, i.e., vulnerability detection. For example, SmartHalo enhances the precision by 21.96% and the recall by 38.00% on detecting integer overflow, improve the precision by 16.67% for contract attack identification. Particularly, we explore the impact of integrating GPT-4o with SmartHalo. With GPT-4o, SmartHalo further enhances its performance, achieving precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
@@
2. Background and Motivation
2.1. Smart Contract Decompiler
@@
Decompilation plays a key role in program analysis by lifting bytecode back into a high-level representation. This process aims to elucidate the underlying logic of contract programs and facilitates a wide range of downstream tasks, such as vulnerability detection (Brent et al., 2020) and component analysis (Wu et al., 2023). Decompilation involves the recovery of variables, functions, and control-flow abstractions through various program analysis methods. Further, it then utilizes heuristic rules to synthesize these abstractions, thereby reconstructing the high-level code representation, commonly referred as decompiled code (Chen et al., 2022a). Recently, an investigation report reveals that over 99% of smart contracts do not make their source code publicly available, with only their bytecode being accessible (Con, [n. d.]). Therefore, smart contract decompilation is becoming increasingly important.
2.2. Deficiencies in Decompiler Output
The current decompiler output exhibits certain deficiencies that obstruct effective comprehension and analysis of the target program. Below we take smart contracts of Fig. 1, which contain source code and corresponding decompilation code, as the instances for illustrating the deficiency of decompiler output.
Deficiency-1. Inaccurate Variable Type Recovery.  Limited by the insufficiently generalized heuristic rule, the SOTA decompilers are inaccurate in recovering the variable type. This deficiency is evident in the example shown in Fig. 1(a). In the source code, the state variable uintStorage is a mapping type that maps bytes32 to uint256 (line 2). However, the decompiler incorrectly assigns it as a mapping uint256 to uint256 (line 10). Additionally, while decompilers narrow down the range of possible types for the parameter varg2 and assign it with bytes (line 11), the correct type of this parameter is actually a string array (line 3). Obviously, the SOTA decompilers commit errors in type inference. The notable efforts in this domain are SigRec (Chen et al., 2022b) and DeepInfer (Zhao et al., 2023), focus primarily on inferring function signatures and recovering parameter types within those signatures for smart contracts. However, since these tools concentrate on a subset of variable types, they are insufficient to address the broader deficiency of type recovery (e.g., variable uintStorage). In addition, the inaccuracies in variable type recovery result in significant difficulty in identifying specific vulnerabilities, such as overflow vulnerabilities (Lai and Luo, 2020).
Deficiency-2. Lack of Contract Attribute.  Due to the limited storage of blockchain, smart contract utilizes state variables (i.e., global variables) to record the key contract attributes (e.g., asset, identity, router). While contract attributes are crucial for the vulnerability analysis task, they are exceptionally difficult to be inferred from decompiled code (Liao et al., 2022). We take an example in Fig. 1(b) for illustration. Contract Bank exhibits a Reentrancy vulnerability because it records the asset balance after transferring assets. An adversary could exploit this by re-entering the contract to transfer assets without recording the change of asset balance (line 7-8). Thus, correctly identifying the state variable Acc.balance as an Asset attribute is critical for detecting this vulnerability. While the contract attribute is explicitly stated in the source code through meaningful variable names (i.e., Acc.balance), it is totally missed at the bytecode level and is labeled as store_0 in its decompiled code (line 11). Hence, inferring the contract attribute for the variable store_0 in the decompiled code becomes a formidable challenge. Without the contract attribute, it is difficult to pinpoint the vulnerability. Prior work (i.e., SmartDagger (Liao et al., 2022)) trains a neural machine learning model using a corpus of 1,200 smart contracts to recover contract attributes for state variables within smart contracts. However, due to limitations in model capacity and the size of the training set, SmartDagger degrade its accuracy of contract attribute prediction when encountering with newly-emerged or rare contracts.
Deficiency-3. Inaccurate Method Identification.  SOTA decompilers are notably ineffective in identifying methods within smart contracts due to their reliance on low-coverage heuristic rules. For instance, the Gigahorse (Grech et al., 2019) decompiler identifies methods by detecting call sites that are invoked recurrently. The heuristic rules result in poor performance when encountering complex methods, particularly inherited methods. In contrast to languages like C++ and Java, where inheritance is clearly delineated, smart contract inheritance involves embedding all base subcontracts ($B_{1},B_{2},...,B_{n}$) as code blocks directly into the inheriting contract $A$, without retaining explicit call information (Sol, [n. d.]). Hence, the heuristic rules failed in this case. As can be seen in Fig. 1(c), the inherited method validate is explicitly stated in the source code (line 6-8), but it is absent at the bytecode level. Our manual investigation reveals that the validate method is integrated into the handle method (lines 10-11). Consequently, SOTA decompilers like Gigahorse are unable to reconstruct such inherited methods, which are critical for downstream program analysis tasks such as cross-contract call flow analysis (Yang et al., 2024; Liao et al., 2022) and component analysis (Sun et al., 2023; Hu et al., 2017).
In conclusion, SOTA decompilers pose significant deficiencies, which have a negative impact on the different kinds of downstream program analysis tasks.
@@ 2.2. Deficiencies in Decompiler Output
Figure 1. Three motivating examples for illustrating the limitations of current decompiler output.
2.3. Our solution
We propose our strategies for optimizing the smart contract decompiler output, combining with a discussion on potential solutions for the deficiencies presented in motivating examples of Fig. 1.
Strategy-1: Using SA to extract control-flow dependency, type dependency and static dependency.  First, type dependency is crucial for the variable type recovery. For example, the type error for variable uintStorage can be corrected by using SA to infer type dependencies, as illustrated in Fig. 1(b). Initially, the output of the predefined function keccak256 is of type bytes32 ($D_{t_{1}}$). Subsequently, the type of variable $v0$ depends on the return value type of the function _getTokenKey, which is also bytes32 ($D_{t_{2}}$). Finally, the key type of the variable setStorage depends on the type of $v0$ ($D_{t_{3}})$. Consequently, the type of setStorage should be corrected to mapping(bytes32= uint256). Second, control-flow dependency is critical for the method boundary identification. For instance, as shown in Fig. 1(a), while static constraints (i.e., call information) are unavailable in method validate, we can manually infer its boundaries by analyzing the contextual control flow, i.e., the method typically declares a new variable at the beginning (line 10) and returns the value at the end (line 11). Third, state dependency is important for contract attribute inference. We take the motivating example in Fig. 1(c) as an instance for illustration. By analyzing the dependency of the state variable, we can find that state variable stor_0 is written in line 16 ($D_{s_{1}}$) and line 18 ($D_{s_{2}}$). By manually analyzing their context usage, we can infer that they are utilized for asset transfer and asset receive, respectively. Hence, we can infer the attribute for stor_0 as the [Asset].
Strategy-2: Empowering LLMs to enrich semantics.  With the remarkable capability of few-shot in-context learning, LLMs are flexible to predicting the optimization target (i.e., variables or methods), which are rarely encountered in the program and often incapable of inferring them based on other variables or methods. We take Fig. 1(b) again as an instance for illustration. The type error for variable input can be corrected by using LLMs for prediction. Fig. 3 presents code optimized by LLM. Benefiting from the extensive training set and model capabilities, LLM can assign the variable input with string memory correctly (line 12). In addition, owing to its remarkable model capacity, LLMs are more effective and generalizable in learning static domain knowledge (i.e., dependencies) for optimization predictions, compared to SA’s manually crafting heuristic rules based on the same knowledge (Peng et al., 2023). For instances, instead of manually summarizing heuristic rules, we can utilize LLMs to learning the context control-flow for method boundary prediction, and to analyze the contextual usage of state variables to infer their contract attributes.
Inspired by the above analysis, our objective is to harness the advantages of both SA and LLMs to optimize decompiler output. Therefore, SmartHalo extracts three types of static dependencies including type dependency, state dependency and control flow dependency within the program through SA. Subsequently, SmartHalo leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these static dependencies.
3. Design of SmartHalo
Figure 2. The decompiled code optimized by LLM for
the instance in Fig. 1(b).

### v2

@@ Abstract
Decompiler is a specialized type of reverse engineering tool extensively employed in program analysis tasks, particularly in program comprehension and vulnerability detection. However, current Solidity smart contract decompilers face significant limitations in reconstructing the original source code. In particular, the bottleneck of SOTA decompilers lies in inaccurate function identification, incorrect variable type recovery, and missing contract attributes. These deficiencies hinder downstream tasks and understanding of the program logic. To address these challenges, we propose SmartHalo, a new framework that enhances decompiler output by combining static analysis (SA) and large language models (LLM). SmartHalo leverages the complementary strengths of SA’s accuracy in control and data flow analysis and LLM’s capability in semantic prediction. More specifically, SmartHalo constructs a new data structure - Dependency Graph (DG), to extract semantic dependencies via static analysis. Then, it takes DG to create prompts for LLM optimization. Finally, the correctness of LLM outputs is validated through symbolic execution and formal verification. Evaluation on a dataset consisting of 465 randomly selected smart contract functions shows that SmartHalo significantly improves the quality of the decompiled code, compared to SOTA decompilers (e.g., Gigahorse). Notably, integrating GPT-4o mini with SmartHalo further enhances its performance, achieving a precision of 91.32% and a recall of 87.38% for function boundaries, a precision of 90.40% and a recall of 88.82% for variable types, and a precision of 80.66% and a recall of 91.78% for contract attributes.
I Introduction
@@
Decompiler is a specific type of reverse engineering tool widely used for program analysis, such as program comprehension and vulnerability detection [1]. For smart contracts written in Solidity, the goal of the decompiler is to recover the machine-executable code (e.g., the EVM bytecode [2]) back to the original source code written by developers (e.g., the Solidity program [3]).
@@
While decompilers recover abstractions that enhance code readability and are widely used by reverse engineers, decompilers are incapable of fully reconstructing the original developer-written code, as the compiler irreversibly subverts crucial information for optimization [4]. In particular, critical information such as variable types, function boundaries, variable names, and annotation are absent in contract bytecode, and are unrecoverable by decompilers. As pointed out by prior research [5, 6, 7, 8], these limitations make the state-of-the-art solidity decompilers, such as Gigahorse [9] and Etherscan decompiler [10], share the following three deficiency in their output: (1) inaccurate function identification: current decompilers often fail to precisely determine function boundaries, leading to the omission of important functions or erroneous function range identification. For example, decompilers often incorrectly recover multiple functions as a single function, which obviously increases complexity in certain tasks (e.g., function-level similarity comparison). (2) inaccurate variable type recovery: decompilers may produce type errors that are inconsistent with static domain rules. For instance, the decompiler often ignores predefined type (e.g., the return value of hash function keccak256 is the type of bytes32), while recovering them as the type of uint256 uniformly. (3) lack of contract attributes. Smart contracts employ state variables to record critical contract attributes (e.g., Asset, identity, router). Although these contract attributes are explicitly stated in the source code through meaningful variable names and annotations, they are entirely omitted in the decompiled code. As highlighted in prior work [5], contract attribute is crucial for vulnerability detection tasks.
@@
Recent advancements [11, 12, 6, 5, 13] start optimizing or reconstructing the absent contract information such as variable type by analyzing the context of decompiler output, even when these information is not part of the contract bytecode. These works are far from satisfied, due to their limited scope and approach-wise weaknesses. More specifically, SmartDagger [5], a framework for detecting cross-contract vulnerability, can partially recover contract attributes (e.g., asset) for state variables from the decompiled bytecode. Neural-FEBI [13] is designed for identifying function boundaries, but it does not support boundary recovery for complex functions such as modifier functions or functions inherited from other contracts/subcontracts. SigRec [12], VarLifter [6] and DeepInfer [11] can partially recover the variable types such as parameter types of already-known function signatures. However, SigRec [12] and VarLifter [6] are heavily relied on pre-defined heuristics according to EVM instructions. Besides, both SmartDagger [5], Neural-FEBI [13] and DeepInfer [11] are based on deep-learning models (e.g., deep neural networks) for information recovery. Therefore, their capability are quite limited to the model-scale and training datasets. For example, they suffers performance degradation when confronted with newly-emerging or rare contracts that are not exist in the training datasets.
@@
Our work.  In this paper, we propose SmartHalo, a novel framework designed to optimize the output of existing smart contract decompilers. In particular, SmartHalo aims at recovering function boundaries, variable types, and contract attributes through a novel combination of static analysis (SA) and large language models (LLMs). The output of SmartHalo can facilitate a number of downstream tasks in program comprehension and security analysis. For example, eliminating false positives and false negatives for vulnerability detection.
@@
Our key insight is that (1) Software exhibits the natural patterns - programmers tend to utilize similar code structures, contract attributes, variable types, and function boundaries in comparable contexts [14]. This repetitiveness enables predicting highly probable contract attributes, function boundaries and variable types for similar contexts. (2) SA and LLMs can collaboratively enhance the output of existing decompilers. Specifically, the advantage of SA lies in its soundness, as it is accurate in handling the optimization targets with complex static constraints [15]. Consider the optimization of variable types as an example. The types of state variables (e.g., address) are inherently related to the program expressions that access or modify them (e.g., msg.sender). In the meantime, the strength of LLM lies in its completeness, as it possesses the flexibility to predict the optimization targets that lack static constraints [16]. For instance, with the remarkable capability of few-shot in-context learning, LLMs are flexible to predicting the types for rare local variables which are incapable of inferring their types based on the types of other variables.
@@
SmartHalo extracts the program dependencies within the program through SA, and then leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these program dependencies. Firstly, SmartHalo extracts three types of dependencies from the decompiled code, including state dependency (e.g., read and write on state variable), control-flow dependency, and type dependency, to construct a fine-grained Dependency Graph (DG) (see Section IV-A). The DG contains static domain knowledge (i.e., dependencies) necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract. Secondly, SmartHalo uses the DG to create prompts that include optimization target contexts (i.e., functions or variables), optimization result candidates (e.g., type or attribute), and chain-of-thoughts that represent inference steps of static analysis for optimization targets. By learning the prompts, LLMs enable the joint optimization on decompiler output with the advantages of both SA and LLMs (see Section IV-B).
@@
A common challenge in LLM-adaptation is to handling its potential hallucination. In our task, LLMs unexpectedly alter the program behaviors of the original decompiled code, and introduce inference errors that contradict common static domain rules (e.g., type knowledge)(see Section IV-C). As countermeasure, SmartHalo conducts the rigorous correctness verification for the LLM output. Specifically, SmartHalo utilizes symbolic execution and formal verification to validate the program-behavior equivalence between the original decompiled code and optimized code. In addition, SmartHalo integrates a set of static rules (e.g., type rules) to identify and reject inference errors.
@@
Evaluation.  To evaluate the effectiveness of SmartHalo, we randomly select 500 functions from the largest open-source contract dataset [17], and manually labelled a dataset containing 456 pairs of source code and decompiler outputs of smart contract functions (44 functions encountered decompilation error). Noting that the dataset size is similar to those used in SOTA studies [4, 18]. The evaluation results indicate that, compared to the original decompiler output (i.e., Gigahorse [9]), SmartHalo (with GPT-3.5) improves the precision and recall of function boundary identification by 20.30% and 30.03%. Further, SmartHalo significantly outperforming tool SOTA VarLifter [6], effectively enhances the precision and recall of variable type inference by 13.51% and 77.08%. Additionally, SmartHalo significantly outperforming SOTA tool SmartDagger [5], successfully improves the precision and recall of contract attributes by 44.69% and 80.86% With the help of SmartHalo, 60.22% of optimized codes can directly be recompiled using the Solidity compiler, whereas the original decompiled output can not support recompilation. We also prove that the optimized output of SmartHalo enhances the effectiveness of downstream tasks, i.e., vulnerability detection. For example, SmartHalo enhances the precision by 21.96% and the recall by 38.00% on detecting integer overflow, improve the precision by 16.67% for contract attack identification.
@@
  • To benefit future research, we release the artifact of SmartHalo, as well as the datasets  [^1] https://github.com/Janelinux/SmartHalo.
II Background and Motivation
II-A Smart Contract Decompiler
Smart contract is a specific type of program running on the blockchain [19], which has enabled a wide range of application scenarios in the digital world [20], such as Decentralized Finance [1], Supply Chain Management [21], and Internet of Things [22]. Recently, the investigation report show that more than 99% smart contracts do not disclose their source code, smart contract decompilation become increasingly important.
@@ II-A Smart Contract Decompiler
Decompilation involves the recovery of variables, functions, and control-flow abstractions through various program analysis methods. Further, it then utilizes heuristic rules to synthesize these abstractions, thereby reconstructing the high-level code representation, commonly referred as decompiled code [16]. Recently, an investigation report reveals that over 99% of smart contracts do not make their source code publicly available, with only their bytecode being accessible [23]. Therefore, smart contract decompilation is becoming increasingly important.
Smart contract decompilation faces significant challenge in accurately restoring bytecode to its original source code [24]. This challenge is primarily caused by two factors: (1) the compiler may lose crucial information about the original program. For example, the lexical parsing stage of the compiler fails to propagate variable types, variable names, annotation and function boundary into the bytecode [25]; and (2) the heuristic rules employed by existing decompilers exhibit low coverage rates [4], This which make the decompiler difficult to adapt to varying contract programming styles or patterns.
@@
Figure 1: Three motivating examples for illustrating the limitations of current decompiler output.
II-B Deficiencies in Decompiler Output
The current decompiler output exhibits certain deficiencies that obstruct effective comprehension and analysis of the target program. Below we take smart contracts of Fig. 1, which contain source code and corresponding decompilation code, as the instances for illustrating the deficiencies.
Deficiency-1. Inaccurate Variable Type Recovery.  Limited by the insufficiently generalized heuristic rule, the SOTA decompilers are inaccurate in recovering the variable type. This deficiency is evident in the example shown in Fig. 1(a). In the source code, the state variable uintStorage is a mapping type that maps bytes32 to uint256 (line 2). However, the decompiler incorrectly assigns it as a mapping uint256 to uint256 (line 10). Additionally, while decompilers narrow down the range of possible types for the parameter varg2 and assign it with bytes (line 11), the correct type of this parameter is actually a string array (line 3). Obviously, the SOTA decompilers commit errors in type inference. The notable efforts in this domain are SigRec [12], VarLifter [6] and DeepInfer [11]. SigRec and DeepInfer focus primarily on inferring function signatures and recovering parameter types within those signatures. Since these tools concentrate on a subset of variable types, they are insufficient to address the broader deficiency of type recovery (e.g., variable uintStorage). VarLifter exhibit low coverage rates, because it heavily relied on pre-defined heuristics according to EVM instructions. In addition, the inaccuracies in variable type recovery result in significant difficulty in identifying specific vulnerabilities, such as overflow vulnerabilities [26].
Deficiency-2. Lack of Contract Attribute.  Due to the limited storage of blockchain, smart contract utilizes state variables (i.e., global variables) to record the key contract attributes (e.g., asset, identity, router). While contract attributes are crucial for the vulnerability analysis task, they are exceptionally difficult to be inferred from decompiled code [5]. We take an example in Fig. 1(b) for illustration. Contract Bank exhibits a Reentrancy vulnerability because it records the asset balance after transferring assets. An adversary could exploit this by re-entering the contract to transfer assets without recording the change of asset balance (line 7-8). Thus, correctly identifying the state variable Acc.balance as an Asset attribute is critical for detecting this vulnerability. While the contract attribute is explicitly stated in the source code through meaningful variable names (i.e., Acc.balance), it is totally missed at the bytecode level and is labeled as store_0 in its decompiled code (line 11). Hence, inferring the contract attribute for the variable store_0 in the decompiled code becomes a formidable challenge. Without the contract attribute, it is difficult to pinpoint the vulnerability. Prior work (i.e., SmartDagger [5]) trains a neural machine learning model using a corpus of 1,200 smart contracts to recover contract attributes for state variables within smart contracts. However, due to limitations in model capacity and the size of the training set, SmartDagger degrade its accuracy of contract attribute prediction when encountering with newly-emerged or rare contracts.
Deficiency-3. Inaccurate function Identification.  SOTA decompilers are notably ineffective in identifying functions within smart contracts due to their reliance on low-coverage heuristic rules. For instance, the Gigahorse [9] decompiler identifies functions by detecting call sites that are invoked recurrently. The heuristic rules result in poor performance when encountering complex functions, particularly inherited functions. In contrast to languages like C++ and Java, where inheritance is clearly delineated, smart contract inheritance involves embedding all base subcontracts ($B_{1},B_{2},...,B_{n}$) as code blocks directly into the inheriting contract $A$, without retaining explicit call information [3]. Hence, the heuristic rules of Gigahorse failed in this case. As can be seen in Fig. 1(c), the inherited function validate is explicitly stated in the source code (line 6-8), but it is absent at the bytecode level. Our manual investigation reveals that the validate function is integrated into the handle function (lines 10-11). Consequently, SOTA decompilers like Gigahorse are unable to reconstruct such inherited functions, which are critical for downstream program analysis tasks such as cross-contract call flow analysis [7, 5] and component analysis [27].
Figure 2: The decompiled code optimized by LLM for the
instance in Fig. 1(b).